OpenStudy (mtalhahassan2):

Need help!! The dephth of water h meters at a seaside resort is given by the formula h= 2.1 sin 2pie(t-3.7)/12.4+5.2. Where t is the time of the day? Use calculater to calculate the time when the water is 7.30 m high.

OpenStudy (anonymous):

can u use the Equation tool? to write it out

OpenStudy (campbell_st):

does this formula give the water level in the resort pool..? so is the equation \[h(t) = 2.1\sin(2\pi \frac{(t - 3.7)}{12.4}) + 5.2\]

OpenStudy (campbell_st):

is t expressed in hours and minutes or in hours so 7:30 is 7.5?